According to a study of U.S. cigarette sales, when the price of cigarettes was 1% higher, consumption would be 0.4% lower in the short run and 0.75% lower in the long run. (Source: Gary Becker, Michael Grossman, and Kevin Murphy, “An Empirical Analysis of Cigarette Addiction”, American Economic Review, Vol. 84 No. 3, June 1994, 396-418.)
a. Calculate the short- and long-run own-price elasticities of the demand for cigarettes.
c. If the government were to impose a tax that raised the price of cigarettes by 5%, what would be the effect on consumer expenditure on cigarettes in the (i) short run, and (ii) long run?
1
Expert's answer
2020-09-21T08:20:20-0400

a)Short-run elasticity is -0,4/1=-0.4; long-run elasticity is -0,75/1=-0.75

c) In case of imposing a tax consumer expenditure will fall by 5×0.4=2% in the short run and 5×0.75=3.75%